# Transient selection of competing order in frustrated spin Peierls systems after a quench

### v1

@@ Abstract
Frustrated magnets represent a prominent class of systems where competing orders appear naturally. Owing to the interplay between the lattice geometry and the exchange interactions, a frustrated magnet hosts distinct magnetic orders that are accidentally degenerate in energy in the mean field limit [4, 5]. Quantum/thermal fluctuations, or chemical/structural disorders may lift this accidental degeneracy through the order by disorder (ObD) mechanism [6, 7, 8], thereby stabilizing one magnetic order out of the many.
Among systems that host competing orders, frustrated magnets are unique in that their free energy landscape are strongly shaped by fluctuations. This trait opens new possibilities for controlling their phases out of the equilibrium. While the much explored Floquet engineering employs the coherent, periodic motion of normal modes [9, 10, 11, 12], less attension is paid to the utility of incoherent fluctuations. In this work, we reveal a mechanism by which transient magnetic orders emerge from such fluctuations in frustrated magnets coupled to soft phonons. After a quench induced by a laser pulse, the system can develop a thermodynamically unstable, competing magnetic order owing to the nonequilibrium phonon fluctuations, before it settles in the equilibrium one.
In equilibrium, a frustrated magnet exhibits the Peierls instability in the presence of spin-phonon coupling  [13, 14, 15, 16]. The instability arises because the system releases the frustration by spontaneously distorting the lattice. When the spin-orbit coupling is weak, the spin-Peierls instability selects a collinear magnetic order among the degenerate magnetic orders.
We examine the evolution of the frustrated spin-Peierls systems after a quench from the paramagnetic state to the magnetically ordered state within the framework of time-dependent Ginzburg-Landau (TDGL) theory [17, 18, 19, 20, 21, 22]. In typical pump-probe experiments on magnets that use femtosecond laser pulses, the magnetic order melts rapidly, and then revives slowly as the system equilibrates [23, 24, 25, 3]. While the microscopic mechanisms for the melting and revival of magnetic order are complex, this process can be phenomenologically modeled by the TDGL formalism as a first approximation.
Figure 1: Transient selection of competing magnetic order in a frustrated spin-Peierls system. (a) Paramagnetic state induced by optical pump. (b) Symmetry equivalent states of the transient non-collinear magnetic order, where the Néel orders in the two sublattices (magenta and cyan arrows) are orthogonal. Black curve sketches the distribution of phonon canonical coordinate $Q$. (c) Thermally stable collinear magnetic order. Exchange interaction are enhanced on green bonds due to spontaneous lattice distortion.
@@
The dynamics of the system falls into the category of the phase ordering process [26, 27, 28]. As both the initial state and the equations of motion preserve all the symmetries, the system develop multiple domains instead of a single phase of the magnetic order in the thermodynamically limit. Particularly, the transient competing order manifests itself as the dominant correlations, rather than true long range order, in the early time, and survive as a subdominant short range order in the late time.
@@
To set the stage, we consider an easy-plane antiferromagnet on the body-centered cubic (BCC) lattice with both nearest neighbor ($J_{1}$) and second neighbor ($J_{2}$) exchange interactions [29]. The BCC lattice comprises of two interpenetrating simple cubic sublattices, dubbed A and B, respectively. When $J_{2}>2J_{1}/3$, each sublattice hosts a Néel order, which are decoupled in the mean field limit (Fig. 1b&c).
Following a standard procedure [30], we write down the symmetry constrained Landau free energy, $F=F_{\mathrm{ex}}+F_{\mathrm{s-ph}}$. $F_{\mathrm{ex}}$ describes the exchange interaction:
@@
$\theta_{A}$ and $\theta_{B}$ describes the Néel order in sublattice A and B, respectively. $K,\kappa$ are spin stiffness. It possesses two $U(1)$ symmetries: The simultaneous rotation of $\theta_{A}$ and $\theta_{B}$ corresponds to the spin rotational symmetry. By contrast, the relative rotational symmetry reflects the accidental degeneracy due to the frustration.
@@
Here, $Q$ is the normal coordinate of an Einstein phonon. The spin-phonon coupling constant $\lambda>0$. Microscopically, it describes the modulation of the $J_{1}$ interactions by lattice distortion (Fig. 1c). $\epsilon$ is the elastic constant of phonons.
@@
$\Gamma_{\mathrm{s}}$ and $\Gamma_{\mathrm{ph}}$ are respectively the kinetic coefficients of the spins and the phonons. $\xi_{A,B,Q}$ are three independent Gaussian white noises. Their second moments are determined by the fluctuation-dissipation theorem (FDT), $\langle\xi_{A,B}(x,t)\xi_{A,B}(x^{\prime},t^{\prime})\rangle=2\Gamma_{\mathrm{s}}k_{B}T\delta^{(3)}(x-y)\delta(t-t^{\prime})$, and $\langle\xi_{Q}(x,t)\xi_{Q}(x^{\prime},t^{\prime})\rangle=2\Gamma_{\mathrm{ph}}k_{B}T\delta^{(3)}(x-y)\delta(t-t^{\prime})$, where $T$ is the bath temperature.
@@
Before embarking on the analysis of the above model, we make a couple of extra steps to simplify it further. We define a pair of fields, $\psi=(\theta_{A}+\theta_{B})/2$, and $\theta=\theta_{A}-\theta_{B}$. The equation of motion for $\psi$ and $\theta$ are then decoupled. $\psi$ is the order parameter corresponding to the spontaneous breaking of the spin rotational symmetry. $\theta$ parametrizes the continuous family of degenerate magnetic orders; the two Néel vectors are collinear when $\theta=0$ or $\pi$, whereas they are orthogonal when $\theta=\pm\pi/2$. Since the focus is the selection of magnetic orders, we drop $\psi$ from now on.
@@
We discretize the model on a cubic lattice for numerical simulation. After appropriate rescaling, we obtain:
$\dot{\theta}_{i}=-\frac{\partial F}{\partial\theta_{i}}+\nu_{i};\quad\tau\dot{Q}_{i}=-\frac{\partial F}{\partial Q_{i}}+\eta_{i},$ (3a)
with the free energy function:
$F=-\sum_{\langle ij\rangle}\cos(\theta_{i}-\theta_{j})-g\sum_{i}Q_{i}\cos\theta_{i}+\sum_{i}\frac{Q^{2}_{i}}{2}.$ (3b)
$g=\sqrt{a^{2}/\epsilon K_{\theta}}\lambda$ is the dimensionless spin-phonon coupling constant. $a$ is lattice cutoff. $K_{\theta}=K(1-\kappa)/2$. $\tau=\Gamma_{s}K_{\theta}/(a^{2}\Gamma_{\mathrm{ph}}\epsilon)$ is the ratio between the characteristic relaxation time of the spins to that of the phonons. Crucially, $\tau\gg 1$ when the phonons are soft ($\epsilon\to 0$). The Gaussian white noises $\nu_{i}$ and $\eta_{i}$ obey the FDT: $\langle\nu_{i}(t)\nu_{j}(t^{\prime})\rangle=2\widetilde{T}\delta_{ij}\delta(t-t^{\prime})$, and $\langle\eta_{i}(t)\eta_{j}(t^{\prime}))\rangle=2\widetilde{T}/\tau\delta_{ij}\delta(t-t^{\prime})$. $\widetilde{T}$ is the dimensionless temperature.
Eq. (3) is the starting point of the ensuing analysis. We set the initial condition to be an uncorrelated state, where $\theta_{i}$ are drawn independently from a uniform distribution over $[0,2\pi)$, and $Q_{i}$ from a Gaussian distribution with variance $\langle Q^{2}_{i}\rangle=3$.
We begin with a qualitative analysis of Eq. (3) in a couple of limits. In the limit of $\tau\to 0$, the phonons response immediately to the spins; therefore, we may set $Q$ to its instantaneous equilibrium value, $\overline{Q}_{i}=g\cos\theta_{i}$. Substituting $Q_{i}$ by $\overline{Q}_{i}$ in the equation of motion for $\theta$, we obtain an effective Landau energy for the spins, $F=-\sum_{\langle ij\rangle}\cos(\theta_{i}-\theta_{j})-g^{2}/2\sum_{i}\cos^{2}\theta_{i}$. The effective anisotropy term generated by the phonons lifts the accidental $U(1)$ degeneracy, and favors the collinear magnetic orders, namely $\theta=0$ or $\pi$. Therefore, the system exhibits the equilibrium magnetic order after the quench.
In the opposite limit of $\tau\to\infty$, $Q$ are frozen to their initial states $Q^{(0)}$. The frozen phonons give rise to a random magnetic field in the $x$ direction, which favors the non-collinear states ($\theta=\pm\pi/2$[32, 33, 34]. To see this, we consider a single domain of size $\xi\gg 1$. Within the domain, the Landau energy has relaxed to its minimum. We write $\theta_{i}=\theta+\delta\theta_{i}$, where $\theta$ is the average orientation of the domain, and $\delta\theta_{i}$ is the local variation due to random field. When $g\ll 1$, expanding $F$ to the second order in $\delta\theta_{i}$ yields $F=1/2\sum_{\langle ij\rangle}(\delta\theta_{i}-\delta\theta_{j})^{2}+g\sin\theta\sum_{i}Q^{(0)}_{i}\delta\theta_{i}$. Minimizing $F$ with respect to $\delta\theta_{i}$, and average over $Q^{(0)}$, we obtain $F=-3\alpha\xi^{3}g^{2}\sin^{2}\theta/2$. Here, $\alpha\approx 0.25$ is a numeric constant. Thus, the frozen phonons generate an effective anisotropy for $\theta$, which favors $\theta=\pm\pi/2$, namely the states where the two Néel orders are orthogonal.
Having understood both limits, we consider the case with finite $\tau$. For simplicity, we drop the white noise at the moment. Solving the equation of motion of $Q$, and substitute the solution into that of $\theta$, we obtain:
$\dot{\theta}_{i}=-\sum_{j\in N_{i}}\sin(\theta_{i}-\theta_{j})-gQ^{0}_{i}e^{-t/\tau}\sin\theta_{i}$
@@
The first summation on the right hand side is over all nearest neighbors. The second term is the random field $\parallel x$ due to the initial fluctuations. According to the preceding discussion, it generates an anisotropy that favors the non-collinear magnetic order. This anisotropy disappears after $t>\tau$. Meanwhile, the third term is a retarded self-interaction mediated by phonons. For the dynamics over time scales $\gg\tau$, the Markov approximation reduces it to $-g^{2}\sin\theta(t)\cos\theta(t)$. This term selects the collinear order similar to the $\tau\to 0$ limit. We thus conclude that the non-collinear magnetic order develops over the time window $t<\tau$, and gives way to the collinear magnetic order when $t>\tau$.
@@
Figure 2: (a) The local spin anisotropy, $\langle\cos(2\theta)\rangle$, as a function of time $t$ for various phonon relaxation times $\tau$. The system size $L=120$. Increasing $L$ doesn’t yield discernible changes. (b) Early time behavior of the local spin anisotropy. (c) Linear scaling between the spin reorientation time $t_{c}$ and the phonon relaxation time $\tau$.
@@
We put this picture to test by numerically solving Eq. (3). We use a $L\times L\times L$ lattice subject to periodic boundary conditions, with the maximal size $L=120$. We choose the representative model parameters $g=0.2$ and $\widetilde{T}=0.1$. The results do not show qualitative change so long as $\lambda\ll 1$ and $\widetilde{T}$ is far below the ordering temperature. We solve the Langevin equation using the Euler method with time step $dt=10^{-2}$. All results are obtained by averaging over 840 runs.
@@
We characterize the magnetic order selection by the observable, $\langle\cos(2\theta)\rangle=\sum_{i}\cos(2\theta_{i})/L^{3}$. $\langle\cos(2\theta)\rangle=1$ if $\theta_{i}=0,\pi$ $\forall i$, and $-1$ if $\theta_{i}=\pm\pi/2$. Thus, this quantity measures the local anisotropy. Fig. 2a shows the numerically calculated $\langle\cos(2\theta)\rangle$ as a function of time. For $\tau=1$, it starts from 0 and increases in time, reflecting the building up of the local collinear magnetic order. By contrast, for $\tau=2000$, $\langle\cos(2\theta)\rangle$ is negative within the simulated time window, indicating the selection of the local non-collinear magnetic order. Its magnitude reaches the maximum at $t\sim 300$. The non-monotonic behavior is due to the fact that the spin anisotropy generated by the phonon fluctuations is slowly decreasing in time.
@@
The threshold stems from the fact that it takes a finite amount of time for the ObD phenomenon to establish. Even for $\tau\to\infty$, $\langle\cos 2\theta\rangle$ starts out being positive, and then becomes negative at $t\approx 2$. We heuristically understand this behavior as follows. The initial dynamics are dominated by the random field $\parallel x$, which gives rise to a small but positive $\langle\cos 2\theta\rangle$. The selection effect appears only after the exchange interaction generates spin correlations across a few lattice spacings. We stress that this behavior is a feature of the ObD physics; the sign change doesn’t occur if the random field is replaced by an anisotropy term $-|\Delta|\sin^{2}\theta$ (Fig. 2b).
@@
Figure 3: (a) Structure factor $S^{x}(q=0)$ as a function of time for various system sizes. Phonon relaxation time $\tau=150$. The inset shows the same data on logarithmic scale. The solid line denotes the $t^{3/2}$ scaling. (b) $S^{y}(q=0)$ as a function of time for various system sizes. The inset shows a comparison between $S^{x}$ and $S^{y}$ for $L=120$ at early times. (c)(d) Correlation length $\xi^{x}$ and $\xi^{y}$ for the collinear and non-collinear magnetic order. The solid line denotes the $t^{1/2}$ scaling.
The observable $\langle\cos 2\theta\rangle$ is designed to probe local magnetic order. We now investigate the non-local spin correlations by measuring the time-dependent structure factor, $S^{a}(\mathbf{q},t)=\sum_{ij}\langle n^{a}_{i}(t)n^{a}_{j}(t)\rangle e^{-i\mathbf{q}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}/L^{3}$. $a=x,y$. $n^{x}_{i}=\cos\theta_{i}$, and $n^{y}_{i}=\sin\theta_{i}$. $S^{x}$ ($S^{y}$) probes the spatial correlation of the (non-)collinear magnetic order.
Fig. 3a shows $S^{x}(q=0)$, the height of the Bragg peak for the collinear magnetic order, as a function time. For prototypical phase ordering process such as that of the Ising model, the Bragg peak height exhibits a $t^{3/2}$ law, reflecting the domain growth [26]. The domain size grows as $\xi\sim t^{1/2}$. As the height is proportional to the domain volume, it grows as $\xi^{3}\sim t^{3/2}$. Here, $S^{x}(q=0)$ exhibits an approximate $t^{3/2}$ law (Fig. 3a, inset).
We extract a correlation length $\xi^{x}$ from $S^{x}(\mathbf{q})$ by measuring the half width at half maximum $q_{0}$ along the $(100)$ direction and using $\xi^{x}=2\pi/q_{0}$ [30]. We observe a deviation from the $t^{1/2}$ growth law. At late time, the dynamics of the model is mapped to a three-dimensional XY model with easy axis anisotropy, which is in the same universality class as the Ising model. The deviation is likely due to uncertainties in measuring the correlation length at late time [30].
@@
For finite system size, $S^{y}(q=0)$ decreases to the equilibrium value $\widetilde{T}/g^{2}\sim O(10)$ at late time because the system is taken over by a single domain of the collinear magnetic order. This is not the case in the thermodynamic limit. As $L$ increases, the slope of the decrease is suppressed. The data therefore suggest $S^{y}(q=0)$ saturates at late time. The anomalously large non-collinear spin correlation is the relic of the transient selection of the competing order.
The fact that the non-collinear magnetic order survives as a subdominant short-range order is mirrored in the correlation length (Fig. 3d). After an initial approximate $t^{1/2}$ growth, $\xi^{y}$ exhibits a kink near the spin reorientation time $t_{c}$ and continues to grow with much smaller slope.

### v2

@@ Abstract
Frustrated magnets represent a prominent class of systems where competing orders appear naturally. A frustrated magnet hosts distinct magnetic orders that are accidentally degenerate in energy [4, 5]. Quantum/thermal fluctuations, or chemical/structural disorders may lift this accidental degeneracy through the order by disorder (ObD) mechanism [6, 7, 8], thereby stabilizing one magnetic order out of the many.
Among systems that host competing orders, frustrated magnets are unique in that their free energy landscape is shaped by fluctuations. This trait opens new possibilities for controlling their phases out of the equilibrium. While the much explored Floquet engineering employs periodic motion of normal modes [9, 10, 11, 12, 13], less attention is paid to incoherent fluctuations. In this work, we reveal a mechanism by which transient magnetic orders emerge from such fluctuations in frustrated magnets coupled to soft phonons. After a quench from the paramagnetic state to magnetically ordered state, the system can develop a thermodynamically unstable, competing magnetic order owing to the nonequilibrium phonon fluctuations, before it settles in the equilibrium one.
In equilibrium, a frustrated magnet exhibits the Peierls instability in the presence of spin-phonon coupling  [14, 15, 16, 17]. The instability arises because the system can release the frustration by spontaneously distorting the lattice. When the spin-orbit coupling is weak, the spin-Peierls instability selects a collinear magnetic order among the degenerate ones.
We examine the evolution of the frustrated spin-Peierls systems after a quench from paramagnetic state to the magnetically ordered state. In typical pump-probe experiments on magnets with femtosecond laser pulses, the magnetic order melts rapidly, and then revives slowly as the system equilibrates [18, 19, 20, 3]. We assume the optical pumping creates a paramagnetic state with an effective bath temperature $T_{i}$. When the pumping is off, the bath cools to a lower temperature $T_{f}$ over a relatively short time scale. We study the revival of the magnetic order as it adapts to the new bath temperature within the framework of time-dependent Ginzburg-Landau (TDGL) theory [21, 22, 23, 24, 25, 26], which exposes the essential ingredients for the physics discussed here. The microscopic mechanisms for the melting and revival of magnetic order are complex. It can nonetheless be phenomenologically modeled by the TDGL formalism as a first approximation. Our protocol is also equivalent to a temperature quench [27, 28].
Figure 1: Transient selection of competing magnetic order in a frustrated spin-Peierls system. (a) Paramagnetic state induced by optical pump. (b) Symmetry equivalent states of the transient non-collinear magnetic order, where the Néel orders in the two sublattices (magenta and cyan arrows) are orthogonal. Black curve sketches the distribution of phonon canonical coordinate $Q$. (c) Thermally stable collinear magnetic order. Exchange interaction are enhanced on green bonds due to spontaneous lattice distortion. The double peaked distribution of $Q$ reflects the two equilibrium positions.
@@
The dynamics of the system falls into the category of the phase ordering process [29, 30, 31]. As the initial state is symmetric, the system develop multiple domains in the thermodynamically limit. The transient competing order manifests itself as the dominant correlations, rather than true long range order, in the early time, and survive as a subdominant short range order in the late time.
@@
We consider an easy-plane, body-centered cubic (BCC) lattice antiferromagnet with both first neighbor ($J_{1}$) and second neighbor ($J_{2}$) exchange interactions,whose ObD physics is among the simplest [7, 32, 33, 34]. The BCC lattice comprises of two interpenetrating simple cubic sublattices, dubbed A and B, respectively. When $J_{2}>2J_{1}/3$, each sublattice hosts a Néel order, which are decoupled in the mean field limit (Fig. 1b&c).
We write down the symmetry constrained Landau free energy, $F=F_{\mathrm{ex}}+F_{\mathrm{s-ph}}$ [35]. $F_{\mathrm{ex}}$ describes the exchange interaction:
@@
$\theta_{A}$ and $\theta_{B}$ describes the direction of Néel order in sublattice A and B, respectively. $K,\kappa$ are spin stiffness. It possesses two $U(1)$ symmetries: The common rotation of $\theta_{A}$ and $\theta_{B}$ corresponds to the spin rotational symmetry. By contrast, the relative rotation reflects the accidental degeneracy due to the frustration. We have used the simplifying approximation that the magnitude of the Néel order parameters are fixed, which means the short-range order is not affected by the quench. This approximation can be removed by using the usual TDGL model, which yields qualitatively similar results [35].
@@
$Q$ is the normal coordinate of an Einstein phonon. The spin-phonon coupling constant $\lambda>0$. Microscopically, it describes the modulation of the $J_{1}$ interactions by the displacement in the non-magnetic ions (not in BCC lattice) that mediate the superexchange (Fig. 1c). $\epsilon>0$ is the elastic constant.
@@
$\Gamma_{\mathrm{s}}$ and $\Gamma_{\mathrm{ph}}$ are respectively the kinetic coefficients of the spins and the phonons. $\xi_{A,B,Q}$ are independent Gaussian white noise fields, whose covariance are $\langle\xi_{\alpha}(x,t)\xi_{\alpha}(x^{\prime},t^{\prime})\rangle=2\Gamma_{\alpha}k_{B}T\delta^{(3)}(x-y)\delta(t-t^{\prime})$. $\Gamma_{\alpha}=\Gamma_{\mathrm{s}}(\Gamma_{\mathrm{Q}})$ when $\alpha=A,B(Q)$. $T$ is the bath temperature.
@@
We make a couple of extra steps to simplify the above model further. We define a pair of fields, $\psi=(\theta_{A}+\theta_{B})/2$, and $\theta=\theta_{A}-\theta_{B}$. The equation of motion for $\psi$ and $\theta$ are then decoupled. $\psi$ is the order parameter corresponding to the spontaneous breaking of the spin rotational symmetry. $\theta$ parametrizes the continuous family of degenerate magnetic orders; the two Néel vectors are collinear when $\theta=0$ or $\pi$, whereas they are orthogonal when $\theta=\pm\pi/2$. Since the focus is the selection of magnetic orders, we drop $\psi$ for now. The resulted minimal model is silent on the magnitude of the Néel order. A direct analysis that treats $\psi$ and $\theta$ on equal footing yields similar behavior in terms of magnetic order selection [35].
@@
We discretize the model on a simple cubic lattice for numerical simulation. After rescaling [35]:
$\dot{\theta}_{i}=-\frac{\partial\widetilde{F}}{\partial\theta_{i}}+\nu_{i};\quad\dot{\widetilde{Q}}_{i}=-\frac{1}{\tau}\frac{\partial\widetilde{F}}{\partial\widetilde{Q}_{i}}+\eta_{i},$ (3a)
with the dimensionless free energy function:
$\widetilde{F}=-\sum_{\langle ij\rangle}\cos(\theta_{i}-\theta_{j})-g\sum_{i}\widetilde{Q}_{i}\cos\theta_{i}+\sum_{i}\frac{\widetilde{Q}^{2}_{i}}{2}.$ (3b)
@@
We estimate $g\sim\sqrt{\delta J/J}$, where $J$ is the spin exchange constant and $\delta J$ its modulation in the spin-Peierls phase. Setting $\delta J/J\sim 10^{-2}$ results in $g\sim 0.1$ [37]. $\tau_{\theta}\sim 1$ to 1010ps estimated from spin wave lifetime [38, 39]. Phonon relaxation time varies greatly  [40, 41], yielding a ratio $\tau$ in the range of 1 to hundreds. The effective model parameters may depend on temperature, which can affect quantitative but not qualitative behavior of the model.
Eq. (3) is the starting point of the ensuing analysis. We set the initial temperature $\widetilde{T}_{i}=3$. As $\widetilde{T}_{i}\gg$ the energy scale for spin exchange and spin-phonon coupling, we approximate the initial state by an uncorrelated state to reduce the computational cost, where $\theta_{i}$ are drawn independently from a uniform distribution over $0,2\pi)$, and $\widetilde{Q}_{i}$ from a Gaussian distribution with $\langle\widetilde{Q}^{2}_{i}\rangle=\widetilde{T}_{i}$ as per the equipartition theorem. Simulations with correlated initial state yield quantitatively similar results [35]. Changing $\widetilde{T}_{i}$ doesn’t alter the qualitative behavior provided it is far above the ordering temperature $\approx 2$.
We begin with a qualitative analysis of Eq. (3) in a couple of limits. In the limit of $\tau\to 0$, the phonons response immediately to the spins; therefore, we may set $\widetilde{Q}$ to its instantaneous equilibrium value, $\overline{Q}_{i}=g\cos\theta_{i}$. Substituting $\widetilde{Q}_{i}$ by $\overline{Q}_{i}$ in the equation of motion for $\theta$, we obtain an effective Landau energy for the spins, $\widetilde{F}=-\sum_{\langle ij\rangle}\cos(\theta_{i}-\theta_{j})-g^{2}/2\sum_{i}\cos^{2}\theta_{i}$. The effective anisotropy term generated by the phonons lifts the accidental $U(1)$ degeneracy, and favors the collinear magnetic orders, namely $\theta=0$ or $\pi$. Therefore, the system exhibits the equilibrium magnetic order after the quench.
In the opposite limit of $\tau\to\infty$, $\widetilde{Q}$ are frozen to their initial states $Q^{(0)}$. The frozen phonons give rise to a random magnetic field in the $x$ direction, which favors the non-collinear states ($\theta=\pm\pi/2$[42, 43, 44]. To see this, we consider a single domain of size $\xi\gg 1$. Within the domain, the Landau energy has relaxed to its minimum. We write $\theta_{i}=\theta+\delta\theta_{i}$, where $\theta$ is the average orientation of the domain, and $\delta\theta_{i}$ is the local variation due to random field. When $g\ll 1$, expanding $\widetilde{F}$ to the second order in $\delta\theta_{i}$ yields $\widetilde{F}=1/2\sum_{\langle ij\rangle}(\delta\theta_{i}-\delta\theta_{j})^{2}+g\sin\theta\sum_{i}Q^{(0)}_{i}\delta\theta_{i}$. Minimizing $\widetilde{F}$ with respect to $\delta\theta_{i}$, and average over $Q^{(0)}$, we obtain $\widetilde{F}=-\alpha\widetilde{T}_{i}\xi^{3}g^{2}\sin^{2}\theta/2$. Here, $\alpha\approx 0.25$ is a constant. Thus, the frozen phonons generate an effective anisotropy for $\theta$, which favors $\theta=\pm\pi/2$, namely the states where the two Néel orders are orthogonal.
@@
Having understood both limits, we consider the case with finite $\tau$. We drop the noise for simplicity. Solving the equation of motion of $\widetilde{Q}$, and substitute the solution into that of $\theta$, we obtain $\dot{\theta}_{i}=-\sum_{j\in N_{i}}\sin(\theta_{i}-\theta_{j})-gQ^{0}_{i}e^{-\widetilde{t}/\tau}\sin\theta_{i}-g^{2}/\tau\int^{\widetilde{t}}_{0}e^{-\frac{\widetilde{t}-s}{\tau}}\sin\theta_{i}(\widetilde{t})\cos\theta_{i}(s)ds$. The first summation on the right hand side is over all nearest neighbors. The second term is the random field $\parallel x$ due to the initial fluctuations. According to the preceding discussion, it generates an anisotropy that favors the non-collinear magnetic order. This anisotropy disappears after $\widetilde{t}>\tau$. Meanwhile, the third term is a retarded self-interaction mediated by phonons. For the dynamics over time scales $\gg\tau$, the Markov approximation reduces it to $g^{2}\sin\theta(t)\cos\theta(t)$. This term selects the collinear order similar to the $\tau\to 0$ limit. We conclude that the non-collinear magnetic order develops over the time window $\widetilde{t}<\tau$, and gives way to the collinear one when $\widetilde{t}>\tau$.
@@
Figure 2: (a) The local spin anisotropy, $\langle\cos(2\theta)\rangle$, as a function of time $t$ for various phonon relaxation times $\tau$. The system size $L=120$. Increasing $L$ doesn’t yield discernible changes. (b) Early time behavior of the local spin anisotropy for various values of $\tau$ and a model with added anisotropy $|\Delta|\sin^{2}\theta$. (c) Linear scaling between the spin reorientation time $t_{c}$ and the phonon relaxation time $\tau$.
@@
We put this picture to test by numerically solving Eq. (3). We use a $L\times L\times L$ lattice subject to periodic boundary conditions, with the maximal size $L=120$. We choose the representative model parameters $g=0.2$ and $\widetilde{T}_{f}=0.1$. The results do not show qualitative change so long as $g\ll 1$ and $\widetilde{T}_{f}$ is far below the ordering temperature. We solve the Langevin equation using the Euler method with time step $dt=10^{-2}$. All results are obtained by averaging over 840 runs.
@@
We characterize the magnetic order selection by the observable, $\langle\cos(2\theta)\rangle=\sum_{i}\cos(2\theta_{i})/L^{3}$. $\langle\cos(2\theta)\rangle=1$ if $\theta_{i}=0,\pi$ $\forall i$, and 1-1 if $\theta_{i}=\pm\pi/2$. Thus, this quantity measures the local anisotropy. Fig. 2a shows the numerically calculated $\langle\cos(2\theta)\rangle$ as a function of time. For $\tau=1$, it starts from 0 and increases in time, reflecting the building up of the local collinear magnetic order. By contrast, for $\tau=2000$, $\langle\cos(2\theta)\rangle$ is negative within the simulated time window, indicating the selection of the local non-collinear magnetic order. Its magnitude reaches the maximum at $\widetilde{t}\sim 300$. The non-monotonic behavior is due to the fact that the spin anisotropy generated by the phonon fluctuations is slowly decreasing in time.
@@
The threshold stems from the fact that it takes a finite amount of time for the ObD phenomenon to establish. Even for $\tau\to\infty$, $\langle\cos 2\theta\rangle$ starts out being positive, and then becomes negative at $\widetilde{t}\approx 2$. We heuristically understand this behavior as follows. The initial dynamics are dominated by the random field $\parallel x$, which gives rise to a small but positive $\langle\cos 2\theta\rangle$. The selection effect appears only after the exchange interaction generates spin correlations across a few lattice spacings. We stress that this behavior is a feature of the ObD physics; the sign change doesn’t occur if the random field is replaced by an anisotropy term $|\Delta|\sin^{2}\theta$ (Fig. 2b).
@@
Figure 3: (a) Structure factor $S^{x}(q=0)$ as a function of time for various system sizes. Phonon relaxation time $\tau=150$. The inset shows the same data on logarithmic scale. The solid line denotes the $t^{3/2}$ scaling. (b) $S^{y}(q=0)$ as a function of time for various system sizes. Red arrow marks the equilibrium value. The inset shows a comparison between $S^{x}$ and $S^{y}$ for $L=120$ at early times. (c)(d) Correlation length $\xi^{x}$ and $\xi^{y}$ for the collinear and non-collinear magnetic order. The solid line denotes the $t^{1/2}$ scaling.
The observable $\langle\cos 2\theta\rangle$ is designed to probe local magnetic order. We now investigate the non-local spin correlations by measuring the time-dependent structure factor, $S^{a}(\mathbf{q},\widetilde{t})=\sum_{ij}\langle n^{a}_{i}(\widetilde{t})n^{a}_{j}(\widetilde{t})\rangle e^{-i\mathbf{q}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}/L^{3}$. $a=x,y$. $n^{x}_{i}=\cos\theta_{i}$, and $n^{y}_{i}=\sin\theta_{i}$. $\langle n^{x,y}_{i}\rangle=0$ by symmetry. $S^{x}$ ($S^{y}$) probes the spatial correlation of the (non-)collinear magnetic order.
Fig. 3a shows $S^{x}(q=0)$, the height of the Bragg peak for the collinear magnetic order, as a function time. For prototypical phase ordering process such as that of the Ising model, the Bragg peak height exhibits a $\widetilde{t}^{3/2}$ law, reflecting the domain growth [29]. The domain size grows as $\xi\sim\widetilde{t}^{1/2}$. As the height is proportional to the domain volume, it grows as $\xi^{3}\sim\widetilde{t}^{3/2}$. Here, $S^{x}(q=0)$ exhibits an approximate $\widetilde{t}^{3/2}$ law (Fig. 3a, inset).
We extract a correlation length $\xi^{x}$ from $S^{x}(\mathbf{q})$ by measuring the half width at half maximum $q_{0}$ along the $(100)$ direction and using $\xi^{x}=2\pi/q_{0}$ [35]. We observe a deviation from the $\widetilde{t}^{1/2}$ growth law. At late time, the dynamics of the model is mapped to a three-dimensional XY model with easy axis anisotropy, which is in the same universality class as the Ising model. The deviation is likely due to uncertainties in measuring the correlation length at late time [35].
@@
For small system size, the system is taken over by a single domain of collinear state with $\theta=0$ or $\pi$ at late time. We expect $S^{y}(q=0)$ will eventually decrease to its thermal equilibrium value (Fig. 3b, red arrow). As $L$ increases, the slope of decrease is suppressed, suggesting it plateaus within numerically accessible time. The large non-collinear spin correlation compared to equilibrium is the relic of the transient selection of the competing order: The late time spin configurations comprise of many domains of collinear states. The non-collinear states survive near the domain walls [35]. It takes exceedingly long time for the non-collinear state to disappear due to the slow motion of the domain walls.
The fact that the non-collinear magnetic order survives as a subdominant short-range order is mirrored in the correlation length (Fig. 3d). After an initial approximate $\widetilde{t}^{1/2}$ growth, $\xi^{y}$ exhibits a kink near the spin reorientation time $t_{c}$ and continues to grow with much smaller slope.